### v2

@@ 6.1.2 Method Comparison
lean
Table 2: Comparison of five explainability methods for hardware trojan detection. M1: 31-member property ensemble with weighted voting. M2: XGBoost with k-NN case retrieval (97.4% correspondence). M3-M5: Single XGBoost (99.56% accuracy) with post-hoc LIME, SHAP, or gradient explanations. Performance metrics reported with 95% bootstrap CIs.
@@
Accuracy (%) 84.45 99.56 99.56 99.56 99.56
Precision (%)b 2.00 46.15 [42.06, 50.24] 46.15 [42.06, 50.24] 46.15 [42.06, 50.24] 46.15 [42.06, 50.24]
Recall (%)b 78.26 52.17 [45.65, 58.70] 52.17 [45.65, 58.70] 52.17 [45.65, 58.70] 52.17 [45.65, 58.70]
F1 Scorec 0.0391 0.4898 0.4898 0.4898 0.4898
@@
Time per sample (ms) <<1.0 <<1.0 33.3 2.07 0.28
[…]
@@
Confidence metric 0.8609 97.4% corresp.
@@
Top Feature LGFi-FFi-PO LGFi (62.8%) PO (34.5%) PO (5.01) LGFi (62.83%)
@@
M1 (property ensemble) trains 31 separate XGBoost classifiers, each on a different combination of the 5 circuit features, then aggregates predictions through weighted voting. This ensemble achieved 84.45% accuracy with average voting confidence of 0.8609 and 100% explanation coverage. However, M1’s precision (2.00%) underperforms M2-M5 (46.15%), reflecting a fundamental limitation: with only five features, devising highly discriminative explainable properties through simple feature subset enumeration is challenging. The 31 feature combinations (all possible subsets of LGFi, FFi, FFo, PI, PO) provide limited semantic meaning compared to domain-aware transformations incorporating circuit-level relationships. While the explanations leverage domain expertise, feature combinations (e.g., LGFi-FFi-PO) align with hardware trojan design principles that security engineers routinely apply during manual circuit analysis. However, the voting mechanism’s aggregation across many weakly discriminative property classifiers produces high recall (78.26%) but low precision. This negative result demonstrates that explainable property-based methods require richer feature engineering beyond naive combinatorics, particularly in low-dimensional settings. The finding is consistent with prior work [32] acknowledging that “it was challenging to map the notion of an explainable property to the minimal feature set” and that “this approach had limited explainability.”
@@
M2 (case-based k-NN) uses a hybrid architecture: XGBoost makes the classification decision, then the system retrieves k=5 nearest neighbors from the training set to provide case-based explanations. This achieves 97.4% average correspondence between the XGBoost prediction and the distance-weighted k-NN vote, demonstrating strong alignment between the classifier and explanatory precedents. Unlike M1’s domain-specific property combinations, M2 provides intuitive explanations through concrete training exemplars. Retrieved neighbors include complete netlist provenance (circuit name, version, line number, net name), enabling engineers to inspect actual training examples to verify structural similarities, a form of explanation that requires minimal domain-specific interpretation.
@@
M2-M5 employ identical XGBoost classification architectures (99.56% accuracy), meaning they produce identical predictions on the test set. The methods differ only in their post-hoc explainability approaches: M2 retrieves k-nearest neighbor training exemplars for case-based explanations, M3 (LIME) uses perturbation-based local linear approximations, M4 (SHAP) employs game-theoretic Shapley values, and M5 (gradient attribution) computes numerical gradients for feature importance. This architectural equivalence explains why M2-M5 show identical confusion matrices in Table 1: the explainability method does not affect classification decisions, only the interpretations provided to users.
@@
M3 (LIME), M4 (SHAP), and M5 (Gradient) all achieved 100% coverage on the test set, generating explanations for all 11,392 samples including all 46 trojan instances. M5 demonstrated the fastest generation time at 0.28ms per sample (119× faster than LIME), while M4 at 2.07ms per sample provided a 16.1× speed advantage over M3’s 33.3ms per sample.
@@ 6.2 Statistical Significance
Bootstrap confidence intervals (10,000 iterations, seed=42) confirmed the statistical robustness of the optimized XGBoost method. At the baseline threshold of 0.5, the precision confidence interval was [7.14%, 13.53%], demonstrating the instability of unoptimized classification. The optimized threshold (0.99) narrowed this interval to [42.06%, 50.24%], indicating more reliable performance.
A McNemar test comparing Method 1 (ensemble) to Method 2 (single XGBoost) found statistically significant differences with $\chi^{2}=1431.25$, p < 0.001. The large chi-square value is partly influenced by the large sample size (n=11,392n=11{,}392). To assess practical significance beyond statistical significance, the effect size is computed: with the observed discordant pairs, the proportions of correct classifications differ between methods (Method 2: 99.56% accuracy vs. Method 1: 84.45%), representing a 15.11 percentage point improvement. Note that Methods 2-5 share identical predictions (same classifier, different explanations), so statistical tests between these methods measure only differences in explainability coverage and computational efficiency, not classification accuracy.
@@ 6.3.1 Feature Importance Rankings
LIME and SHAP methods revealed complementary feature importance patterns for trojan detection. LIME’s top-ranked features (by frequency of #1 ranking) were: PO (34.5%), FFo (17.5%), LGFi (17.3%), FFi (2.0%), and PI (1.9%). SHAP’s top features (by average absolute SHAP value) were: PO (5.0137), LGFi (4.7780), FFo (4.2500), PI (1.9937), and FFi (1.5458).
@@ 6.3.2 Explanation Coverage and Computational Efficiency
SHAP’s TreeExplainer implementation exploited the tree structure of XGBoost models to achieve 16.1-fold speedup over LIME’s perturbation-based approach. For the 11,392-sample test set, SHAP required approximately 23.5 seconds total explanation time versus LIME’s 379 seconds, making SHAP more suitable for production deployment scenarios requiring real-time or near-real-time explanation generation.
@@ 6.4 Discussion
Optimized threshold selection (0.99) improves trojan detection performance, transforming XGBoost from a high-recall but low-precision classifier (baseline: 76.09% recall, 10.17% precision) to a more balanced detector (52.17% recall, 46.15% precision). This threshold optimization trades some recall for a 4.5-fold precision improvement, reducing false positive rates from 2.66% to 0.25%.
The complementarity of property-based and case-based explainability approaches suggests that hybrid systems could combine the inherent interpretability of property ensembles for high-coverage explanations with model-agnostic methods to explain higher-performing but opaque models. The 16.1-fold speed advantage of SHAP over LIME makes it the preferred model-agnostic explanation method for practical trojan detection workflows.
Comparison with the Hasegawa et al. [4] baseline (5.13% precision at 5.6% FPR) demonstrates that modern gradient-boosted methods with careful threshold optimization achieve order-of-magnitude improvements in precision (9-fold) and recall (37-fold) while maintaining lower false positive rates (0.25% vs 5.6%).
@@ 6.4.1 Property-Based Explainability in Low-Dimensional Settings
The property-based method’s underperformance (M1: 2.00% precision vs. M2-M5: 46.15% precision) reflects a fundamental limitation of domain-aware explainability in feature-constrained domains. With only five structural features extracted from gate-level netlists, the 31 property ensemble, comprising all possible feature subsets, lacks the discriminative power and semantic richness needed for effective trojan detection. This negative result demonstrates that explainable AI architectures cannot rely solely on combinatorial enumeration of minimal features.
Three directions for improving property-based explainability in low-dimensional settings are possible. First, ratio-based properties could encode domain relationships explicitly: $\text{LGFi}/\text{PO}$ captures “fanin complexity near outputs” (trigger signature), $\text{PI}/\text{PO}$ indicates “input-to-output proximity balance” (trigger vs. payload indicator), and $\text{FFi}/\text{FFo}$ measures “flip-flop isolation asymmetry” (sequential depth patterns). Second, interaction terms mimicking XGBoost’s learned relationships could provide interpretable alternatives: $\text{LGFi}\times\text{FFo}$ represents “complex logic with downstream propagation paths” (payload characteristic), while $\text{PI}\times\text{LGFi}$ captures “input proximity with fanin complexity” (trigger activation patterns). Third, circuit-relative percentile features address the rare-event nature of trojans: instead of absolute feature values that vary by circuit size, percentile ranks (e.g., 95th percentile LGFi within circuit) identify anomalous gates relative to their circuit context, improving cross-circuit generalization.
These feature engineering approaches align with the observation [32] that property-based methods work best for high-dimensional data where rich feature spaces enable diverse property discovery, while case-based methods excel in low-dimensional settings. The current work’s comparison validates this tradeoff: M2’s case-based hybrid achieves both high performance (99.56% accuracy, 46.15% precision) and strong explainability (97.4% correspondence), suggesting that for minimal feature sets, combining domain-agnostic ML models with precedent-based explanations may be more effective than property ensemble voting. However, semantically meaningful feature transformations (ratios, interactions, percentiles) are underexplored for domain-aware explainability in constrained settings.
@@
[…]
The speed difference reflects fundamental algorithmic differences. Gradient attribution requires only 6 model evaluations per sample (baseline plus one perturbation per feature), computing numerical derivatives via finite differences with $\epsilon=0.01$. SHAP’s TreeExplainer, while optimized for tree ensembles through polynomial-time tree traversal, still requires analyzing conditional expectations across feature subsets to compute exact Shapley values. For the 11,392-gate test set, gradient attribution completes in approximately 3.2 seconds versus SHAP’s 23.5 seconds, a difference that compounds in production deployments processing millions of gates.
However, computational efficiency trades theoretical rigor for speed. SHAP provides formal guarantees of local accuracy, missingness, and consistency: the sum of feature attributions exactly equals the difference between prediction and baseline, and attributions remain consistent across models. Gradient attribution lacks these guarantees: finite differences provide only local linear approximations that may be inaccurate for nonlinear decision boundaries. For XGBoost tree ensembles, this limitation is mitigated because tree-based models are piecewise constant, making gradients well-defined within each leaf region. Nevertheless, gradient attributions represent approximations rather than exact decompositions.
@@
[…]
@@
For hardware trojan detection specifically, the domain-opaque nature of both gradient and SHAP attributions limits their practical utility. While gradient’s speed advantage matters for large-scale deployment, these feature attribution methods lack the circuit-level context (“high fanin near outputs indicates trigger circuits”) needed for validation and remediation. The comparison reveals a tradeoff: domain-aware property analysis aligns with engineering expertise, model-agnostic case-based reasoning provides precedent without domain requirements, while feature attribution methods offer mathematical rigor at the cost of domain context. This finding has implications beyond hardware security for domains requiring interpretable ML predictions.
@@ 7 Conclusion
This work provides empirical evidence comparing explainability approaches for hardware trojan detection across three categories: domain-aware property analysis, model-agnostic case-based reasoning, and model-agnostic feature attribution. Through systematic comparison of five methods on gate-level netlist analysis, the findings show distinct tradeoffs. Domain-aware property analysis provides circuit-specific interpretations aligned with engineering expertise. Case-based reasoning offers model-agnostic explanations through training exemplars that require minimal domain interpretation. Feature attribution methods (LIME, SHAP, Gradient) provide mathematically rigorous but domain-opaque importance scores.
@@
The comparative evaluation on 11,392 gates from the Trust-Hub benchmark dataset yields several results. For detection performance, XGBoost classification with optimized decision threshold achieves 46.15% precision (95% CI: [42.06%, 50.24%]) and 52.17% recall (95% CI: [45.65%, 58.70%]) on severely imbalanced data (275:1 benign-to-trojan ratio), representing a 9-fold improvement in precision over prior work (Hasegawa et al.: 5.13% precision, 82.6% recall). This improvement reduces false positive rates from approximately 5.6% to 0.25%, addressing a limitation for practical deployment where security engineers must manually investigate flagged gates. Statistical validation via McNemar’s test confirms these improvements are statistically significant ($\chi^{2}=1431.25$, $p<0.001$), with the 15.11 percentage point accuracy improvement (99.56% vs. 84.45%) indicating meaningful practical differences beyond statistical significance alone.
@@
For explainability characteristics, property-based analysis provides domain-aligned interpretations through 31 circuit-specific patterns like "high fanin complexity (LGFi=12) near primary output (PO=1) indicates potential rare-event trigger circuits." Case-based reasoning achieves 97.4% correspondence between predictions and nearest training neighbors, providing model-agnostic justifications grounded in precedent with complete provenance to enable verification. LIME and SHAP produce mathematically rigorous feature attributions with strong inter-method correlation ($\rho=0.94$, $p<0.001$), but provide generic importance scores like "FFo importance +0.0526" without circuit-level interpretation. Gradient attribution achieves 481× computational speedup over SHAP (0.52ms vs. 250ms per explanation) with similar generic attributions.
@@
For practitioners selecting explainability methods for hardware trojan detection deployments, the results suggest the following guidance. Property-based analysis suits scenarios where domain-aligned explanations enable security engineers to validate findings through circuit design knowledge and plan targeted remediation. Case-based reasoning suits scenarios where precedent-based justification builds trust, enabling engineers to inspect actual training examples without requiring deep circuit expertise for each explanation. Feature attribution methods (LIME, SHAP) suit exploratory analysis where mathematical rigor and inter-method consistency matter, though generic importance scores require domain expertise to interpret for security decisions. Gradient attribution offers computational efficiency for large-scale screening with similar generic attributions.
@@
The limitations of this study provide context for interpreting results and identifying future research directions. First, evaluation is restricted to the Trust-Hub benchmark dataset with synthetic trojan insertions; generalization beyond this controlled environment requires careful consideration. The evaluation focuses exclusively on digital combinational trojans in gate-level netlists, excluding analog/mixed-signal trojans, sequential trigger mechanisms, and system-level threats. While the dataset includes circuits synthesized with distinct standard cell libraries providing synthesis diversity, all circuits are academic benchmarks limiting conclusions about generalizability to industrial designs. Real-world hardware from untrusted fabrication facilities may exhibit trojan characteristics different from academic benchmarks, requiring validation on actual deployed systems when such datasets become available. Second, the five-feature representation captures structural properties but omits functional characteristics, timing information, and power signatures that may provide complementary detection signals. Third, extreme class imbalance (275:1) poses challenges: while 52.17% recall represents measurable improvement over prior work, 47.83% of trojans remain undetected, requiring complementary detection techniques for comprehensive security coverage. Fourth, the evaluation focuses on quantitative metrics and qualitative inspection rather than controlled human-subjects studies with practicing hardware security engineers, limiting conclusions about real-world interpretability and trust.
@@
Future research directions include: (1) Multi-dataset validation across different circuit families, standard cell libraries, and trojan insertion methods to assess generalizability beyond Trust-Hub; (2) Feature space expansion incorporating functional simulation, timing analysis, and side-channel signatures to improve detection coverage, with specific emphasis on domain-aware feature engineering for property-based explainability: ratio-based properties (e.g., LGFi/PO for “fanin near outputs”), interaction terms (e.g., $\text{LGFi}\times\text{FFo}$ for “complex logic with downstream paths”), and circuit-relative percentile features to capture rare-event signatures while maintaining interpretability; (3) Advanced imbalance handling through synthetic oversampling, cost-sensitive learning, and ensemble methods to address the 275:1 class imbalance; (4) Human-subjects evaluation with practicing security engineers to quantify actionability, trust, and decision-making effectiveness across explainability methods; (5) Hybrid explainability approaches combining domain-aware circuit analysis with model-agnostic attribution to combine strengths of both paradigms; (6) Extension to sequential trojan detection where multiple stages of analysis progressively refine suspicion rankings; (7) Natural language explanation synthesis using large language models [36, 37] to transform structured explainability outputs (property patterns, feature attributions, case exemplars) into reports accessible to verification engineers with varying expertise levels, with human-subjects evaluation to validate usability improvements.
The broader implication for explainable AI deployment extends beyond hardware security. This work demonstrates that in specialized domains where practitioners possess deep domain expertise, domain-aware explainability methods that align with existing reasoning patterns may provide better actionability compared to general-purpose model-agnostic techniques, despite the latter’s mathematical rigor and universal applicability. These findings have relevance for XAI deployment across safety-critical and security-sensitive domains including medical diagnosis, autonomous systems, financial fraud detection, and infrastructure protection, where domain-specific interpretability may be valuable for trust, validation, and regulatory compliance. The tradeoff between domain-agnostic generality and domain-specific actionability represents a consideration for XAI system design in practice.

### v3

@@ 6.1.2 Method Comparison
Table 2: Comparison of five explainability methods for hardware trojan detection. M1: 31-member property ensemble with weighted voting. M2: XGBoost with k-NN case retrieval (95.82% correspondence). M3-M5: Single XGBoost (99.57% accuracy) with post-hoc LIME, SHAP, or gradient explanations. Performance metrics reported with 95% bootstrap CIs.
@@
Accuracy (%) 66.54 99.57 99.57 99.57 99.57
Precision (%)b 1.78 72.06 72.06 72.06 72.06
   95% CI [1.38, 2.21] [61.11, 82.54] [61.11, 82.54] [61.11, 82.54] [61.11, 82.54]
Recall (%)b 87.34 62.03 62.03 62.03 62.03
   95% CI [79.31, 94.20] [51.22, 72.58] [51.22, 72.58] [51.22, 72.58] [51.22, 72.58]
F1 Scorec 0.0349 0.6667 0.6667 0.6667 0.6667
@@
Time per sample (ms) <<1.0 <<1.0 24.4 1.10 0.16
Speedup vs. LIME baseline 22 speedup factor 150 speedup factor
Speedup vs. SHAP 0.045 baseline 7 speedup factor
@@
Confidence metric 0.8244 95.82% corresp.
@@
Top Feature LGFi-FFi-PO LGFi (62.8%) PO (38.5%) PO (4.11) LGFi (62.83%)
@@
M1 (property ensemble) trains 31 separate XGBoost classifiers, each on a different combination of the 5 circuit features, then aggregates predictions through weighted voting. This ensemble achieved 66.54% accuracy with average voting confidence of 0.8244 and 100% explanation coverage. However, M1’s precision (1.78%) underperforms M2-M5 (72.06%), reflecting a fundamental limitation: with only five features, devising highly discriminative explainable properties through simple feature subset enumeration is challenging. The 31 feature combinations (all possible subsets of LGFi, FFi, FFo, PI, PO) provide limited semantic meaning compared to domain-aware transformations incorporating circuit-level relationships. The explanations draw on domain expertise, as feature combinations (e.g., LGFi-FFi-PO) align with hardware trojan design principles that security engineers routinely apply during manual circuit analysis. However, the voting mechanism’s aggregation across many weakly discriminative property classifiers produces high recall (87.34%) but low precision. This negative result demonstrates that explainable property-based methods require richer feature engineering beyond naive combinatorics, particularly in low-dimensional settings. The finding is consistent with prior work [32] acknowledging that “it was challenging to map the notion of an explainable property to the minimal feature set” and that “this approach had limited explainability.”
@@
M2 (case-based k-NN) uses a hybrid architecture: XGBoost makes the classification decision, then the system retrieves k=5 nearest neighbors from the training set to provide case-based explanations. This achieves 95.82% average correspondence between the XGBoost prediction and the distance-weighted k-NN vote, demonstrating strong alignment between the classifier and explanatory precedents. Unlike M1’s domain-specific property combinations, M2 provides intuitive explanations through concrete training exemplars. Retrieved neighbors include complete netlist provenance (circuit name, version, line number, net name), enabling engineers to inspect actual training examples to verify structural similarities, a form of explanation that requires minimal domain-specific interpretation.
@@
M2-M5 employ identical XGBoost classification architectures (99.57% accuracy), meaning they produce identical predictions on the test set. The methods differ only in their post-hoc explainability approaches: M2 retrieves k-nearest neighbor training exemplars for case-based explanations, M3 (LIME) uses perturbation-based local linear approximations, M4 (SHAP) employs game-theoretic Shapley values, and M5 (gradient attribution) computes numerical gradients for feature importance. This architectural equivalence explains why M2-M5 show identical confusion matrices in Table 2: the explainability method does not affect classification decisions, only the interpretations provided to users.
@@
M3 (LIME), M4 (SHAP), and M5 (Gradient) all achieved 100% coverage on the test set, generating explanations for all 11,392 samples including all 79 trojan instances. M5 demonstrated the fastest explanation time at 0.16ms per sample (a speedup factor of 150 over LIME), while M4 at 1.10ms per sample provided a speedup factor of 22 over M3’s 24.4ms per sample.
@@ 6.2 Statistical Significance
Bootstrap confidence intervals (10,000 iterations, seed=42) confirmed the statistical robustness of the optimized XGBoost method. At the baseline threshold of 0.5, the precision confidence interval was [9.21%, 14.56%] and recall was [71.87%, 89.19%], demonstrating the instability of unoptimized classification. At the optimized threshold of 0.980, precision improved to [61.11%, 82.54%] and recall was [51.22%, 72.58%], with the F1 confidence interval of [0.5714, 0.7500] confirming that the performance gain is robust across bootstrap samples.
A McNemar’s chi-squared test comparing Method 1 (ensemble) to Method 2 (single XGBoost) found statistically significant differences ($\chi^{2}(1)=3305.07$, $p<0.001$). The large chi-squared value is partly influenced by the large sample size (n=11,392n=11{,}392). To assess practical significance beyond statistical significance, the effect size is computed: with the observed discordant pairs, the proportions of correct classifications differ between methods (Method 2: 95.69% accuracy at the default threshold vs. Method 1: 66.54%), representing a 29.15 percentage point improvement. Note that Methods 2-5 share identical predictions (same classifier, different explanations), so statistical tests between these methods measure only differences in explainability coverage and computational efficiency, not classification accuracy.
@@ 6.3.1 Feature Importance Rankings
LIME and SHAP methods revealed complementary feature importance patterns for trojan detection. LIME’s top-ranked features (by frequency of #1 ranking) were: PO (38.5%), FFo (26.7%), LGFi (22.3%), PI (10.8%), and FFi (1.7%). SHAP’s top features (by average absolute SHAP value) were: PO (4.1071), LGFi (3.4160), FFo (2.8944), PI (1.9109), and FFi (1.4086).
@@ 6.3.2 Explanation Coverage and Computational Efficiency
SHAP’s TreeExplainer implementation exploited the tree structure of XGBoost models to achieve a 22 speedup factor over LIME’s perturbation-based approach. For the 11,392-sample test set, SHAP required approximately 13 seconds total explanation time versus LIME’s 278 seconds, making SHAP more suitable for production deployment scenarios requiring real-time or near-real-time explanation generation.
@@ 6.4 Discussion
Optimized threshold selection (0.980) improves trojan detection performance, transforming XGBoost from a high-recall but low-precision classifier (baseline: 81.00% recall, 11.85% precision) to a more balanced detector (62.03% recall, 72.06% precision). This threshold optimization trades some recall for a 6.1-fold precision improvement, reducing false positive rates from 4.21% to 0.17%.
The complementarity of property-based and case-based explainability approaches suggests that hybrid systems could combine the inherent interpretability of property ensembles for high-coverage explanations with model-agnostic methods to explain higher-performing but opaque models. The 22 speedup factor of SHAP over LIME makes it the preferred model-agnostic explanation method for practical trojan detection workflows.
Comparison with the Hasegawa et al. [4] baseline (5.13% precision at 50.5% FPR) demonstrates that modern gradient-boosted methods with careful threshold optimization achieve order-of-magnitude improvements in precision (14-fold at the optimized threshold, 2.3-fold at default) while dramatically reducing false positive rates (0.17% vs 50.5%). This work’s recall (62.03%) is lower than Hasegawa’s (82.6%), reflecting the precision–recall tradeoff inherent in threshold optimization. This work retains Hasegawa’s original five-feature representation to enable direct comparison and to maintain interpretable circuit characteristics. Expanding the feature set could improve detection performance but would shift the focus away from the explainability analysis that is this work’s primary contribution.
@@ 6.4.1 Property-Based Explainability in Low-Dimensional Settings
The property-based method’s underperformance (M1: 1.78% precision vs. M2-M5: 72.06% precision) reflects a fundamental limitation of domain-aware explainability in feature-constrained domains. With only five structural features extracted from gate-level netlists, the 31 property ensemble, comprising all possible feature subsets, lacks the discriminative power and semantic richness needed for effective trojan detection. This negative result demonstrates that explainable AI architectures cannot rely solely on combinatorial enumeration of minimal features.
Three directions for improving property-based explainability in low-dimensional settings are possible:
  1. 1.
    Ratio-based properties could encode domain relationships explicitly as quotients of feature pairs: $\text{LGFi}/\text{PO}$ (fanin complexity divided by output distance) captures “fanin complexity near outputs” as a trigger signature; $\text{PI}/\text{PO}$ (input distance divided by output distance) indicates “input-to-output proximity balance” as a trigger vs. payload indicator; and $\text{FFi}/\text{FFo}$ (upstream flip-flop distance divided by downstream flip-flop distance) measures “flip-flop isolation asymmetry” reflecting sequential depth patterns.
  2. 2.
    Interaction terms could mimic XGBoost’s learned feature relationships through pairwise products: the product $\text{LGFi}\cdot\text{FFo}$ (fanin complexity multiplied by downstream flip-flop distance) represents “complex logic with downstream propagation paths” as a payload characteristic, while the product $\text{PI}\cdot\text{LGFi}$ (input proximity multiplied by fanin complexity) captures “input proximity with fanin complexity” as a trigger activation pattern.
  3. 3.
    Circuit-relative percentile features could address the rare-event nature of trojans: instead of absolute feature values that vary by circuit size, percentile ranks (e.g., 95th percentile LGFi within circuit) would identify anomalous gates relative to their circuit context, improving cross-circuit generalization.
These feature engineering approaches align with the observation [32] that property-based methods work best for high-dimensional data where rich feature spaces enable diverse property discovery, while case-based methods excel in low-dimensional settings. The current work’s comparison validates this tradeoff: M2’s case-based hybrid achieves both high performance (99.57% accuracy, 72.06% precision) and strong explainability (95.82% correspondence), suggesting that for minimal feature sets, combining domain-agnostic ML models with precedent-based explanations may be more effective than property ensemble voting. However, semantically meaningful feature transformations (ratios, interactions, percentiles) are underexplored for domain-aware explainability in constrained settings.
@@
The gradient attribution method (M5) achieves 0.16ms per explanation, representing 150 speedup factor over LIME (24.4ms) and 7 speedup factor over SHAP (1.10ms). Despite this speed advantage, gradient and SHAP produce similar feature importance rankings for individual gates, with both methods identifying the same most-influential features in the majority of cases. This raises a question for practitioners: when does computational efficiency justify sacrificing SHAP’s theoretical guarantees?
The speed difference reflects fundamental algorithmic differences. Gradient attribution requires only 6 model evaluations per sample (baseline plus one perturbation per feature), computing numerical derivatives via finite differences with an error $\epsilon=0.01$. SHAP’s TreeExplainer, while optimized for tree ensembles through polynomial-time tree traversal, still requires analyzing conditional expectations across feature subsets to compute exact Shapley values. For the 11,392-gate test set, gradient attribution completes in approximately 1.9 seconds versus SHAP’s 13 seconds, a difference that compounds in production deployments processing millions of gates.
However, computational efficiency trades formal guarantees for speed. SHAP provides formal guarantees of local accuracy, missingness, and consistency: the sum of feature attributions exactly equals the difference between prediction and baseline, and attributions remain consistent across models. Gradient attribution lacks these guarantees: finite differences provide only local linear approximations that may be inaccurate for nonlinear decision boundaries. For XGBoost tree ensembles, this limitation is mitigated because tree-based models are piecewise constant, making gradients well-defined within each leaf region. Nevertheless, gradient attributions represent approximations rather than exact decompositions.
@@
  • Production deployment: Gradient attribution suits real-time systems requiring sub-millisecond explanations for interactive debugging or high-throughput screening. The 7 speedup factor over SHAP enables explaining every gate in large circuits (1M+ gates) within minutes rather than hours. Accept approximate attributions for operational efficiency.
  • Research and validation: SHAP provides publication-quality explanations with formal guarantees for rigorous analysis. The 22 speedup factor over LIME makes SHAP practical for research datasets (10k to 100k samples) while preserving theoretical foundations. Use for peer-reviewed studies and regulatory documentation.
@@
For hardware trojan detection specifically, the domain-opaque nature of both gradient and SHAP attributions limits their practical utility. While gradient’s speed advantage matters for large-scale deployment, these feature attribution methods lack the circuit-level context (“high fanin near outputs indicates trigger circuits”) needed for validation and remediation. The comparison reveals a tradeoff: domain-aware property analysis aligns with engineering expertise, model-agnostic case-based reasoning provides precedent without domain requirements, while feature attribution methods offer formal precision at the cost of domain context. This finding has implications beyond hardware security for domains requiring interpretable ML predictions.
@@ 7 Conclusion
This work provides empirical evidence comparing explainability approaches for hardware trojan detection across three categories: domain-aware property analysis, model-agnostic case-based reasoning, and model-agnostic feature attribution. Through systematic comparison of five methods on gate-level netlist analysis, the findings show distinct tradeoffs. Domain-aware property analysis provides circuit-specific interpretations aligned with engineering expertise. Case-based reasoning offers model-agnostic explanations through training exemplars that require minimal domain interpretation. Feature attribution methods (LIME, SHAP, Gradient) provide mathematically principled but domain-opaque importance scores.
@@
The comparative evaluation on 11,392 gates from the Trust-Hub benchmark dataset yields several results. For detection performance, XGBoost classification with optimized decision threshold achieves 72.06% precision (95% CI: [61.11%, 82.54%]) and 62.03% recall (95% CI: [51.22%, 72.58%]) on severely imbalanced data (158 to one benign-to-trojan ratio), representing a 14-fold improvement in precision at the optimized threshold over prior work (Hasegawa et al.: 5.13% precision, 82.6% recall), or 2.3-fold at the default threshold. This improvement reduces false positive rates from approximately 50.5% to 0.17%, addressing a limitation for practical deployment where security engineers must manually investigate flagged gates. Statistical validation via McNemar’s chi-squared test ($\chi^{2}(1)=3305.07$, $p<0.001$) confirms that the classification differences between Method 1 (property ensemble) and Methods 2–5 (single XGBoost) are statistically significant. The 29.15 percentage point accuracy improvement (95.69% at the default threshold vs. 66.54%) indicates meaningful practical differences beyond statistical significance alone.
@@
For explainability characteristics, property-based analysis provides domain-aligned interpretations through 31 circuit-specific patterns like “high fanin complexity (LGFi=12) near primary output (PO=1) indicates potential rare-event trigger circuits.” Case-based reasoning achieves 95.82% correspondence between predictions and nearest training neighbors, providing model-agnostic justifications grounded in precedent with complete provenance to enable verification. LIME and SHAP produce mathematically rigorous feature attributions with strong inter-method Spearman rank correlation ($\rho_{s}=1.000$, $p<0.001$, where $\rho_{s}$ is the Spearman correlation coefficient and $p$ is the p-value), but provide generic importance scores like “FFo importance +0.0526” without circuit-level interpretation. Gradient attribution achieves a 7 computational speedup factor over SHAP (0.16ms vs. 1.10ms per explanation) with similar generic attributions.
@@
For practitioners selecting explainability methods for hardware trojan detection deployments, the results suggest the following guidance. Property-based analysis suits scenarios where domain-aligned explanations enable security engineers to validate findings through circuit design knowledge and plan targeted remediation. Case-based reasoning suits scenarios where precedent-based justification builds trust, enabling engineers to inspect actual training examples without requiring deep circuit expertise for each explanation. Feature attribution methods (LIME, SHAP) suit exploratory analysis where mathematical techniques and inter-method consistency matter, though generic importance scores require domain expertise to interpret for security decisions. Gradient attribution offers computational efficiency for large-scale screening with similar generic attributions.
@@
The limitations of this study provide context for interpreting results and identifying future research directions. First, evaluation is restricted to the Trust-Hub benchmark dataset with synthetic trojan insertions. Generalization beyond this controlled environment requires careful consideration. The evaluation focuses exclusively on digital combinational trojans in gate-level netlists, excluding analog/mixed-signal trojans, sequential trigger mechanisms, and system-level threats. While the dataset includes circuits synthesized with distinct standard cell libraries providing synthesis diversity, all circuits are academic benchmarks limiting conclusions about generalizability to industrial designs. Real-world hardware from untrusted fabrication facilities may exhibit trojan characteristics different from academic benchmarks, requiring validation on actual deployed systems when such datasets become available. Second, the five-feature representation captures structural properties but omits functional characteristics, timing information, and power signatures that may provide complementary detection signals. Third, extreme class imbalance (143 to one) poses challenges: while 62.03% recall represents measurable improvement over prior work, 37.97% of trojans remain undetected, requiring complementary detection techniques for complete security coverage. Fourth, the evaluation focuses on quantitative metrics and qualitative inspection rather than controlled human-subjects studies with practicing hardware security engineers, limiting conclusions about real-world interpretability and trust.
@@
Future research directions include: (1) Multi-dataset validation across different circuit families, standard cell libraries, and trojan insertion methods to assess generalizability beyond Trust-Hub; (2) Feature space expansion incorporating functional simulation, timing analysis, and side-channel signatures to improve detection coverage, with specific emphasis on domain-aware feature engineering for property-based explainability: ratio-based properties (e.g., LGFi/PO for “fanin near outputs”), interaction terms (e.g., $\text{LGFi}\times\text{FFo}$ for “complex logic with downstream paths”), and circuit-relative percentile features to capture rare-event signatures while maintaining interpretability; (3) Advanced imbalance handling through synthetic oversampling, cost-sensitive learning, and ensemble methods to address the 158 to one class imbalance; (4) Human-subjects evaluation with practicing security engineers to quantify actionability, trust, and decision-making effectiveness across explainability methods; (5) Hybrid explainability approaches combining domain-aware circuit analysis with model-agnostic attribution to combine strengths of both paradigms; (6) Extension to sequential trojan detection where multiple stages of analysis progressively refine suspicion rankings; (7) Natural language explanation synthesis using large language models [41, 42] to transform structured explainability outputs (property patterns, feature attributions, case exemplars) into reports accessible to verification engineers with varying expertise levels, with human-subjects evaluation to validate usability improvements.